# Fine-Grained Retrieval-Augmented Generation for Visual Question Answering

### v1

@@ Abstract
Visual Question Answering (VQA) focuses on providing answers to natural language questions by utilizing information from images. Although cutting-edge multimodal large language models (MLLMs) such as GPT-4o achieve strong performance on VQA tasks, they frequently fall short in accessing domain-specific or the latest knowledge. To mitigate this issue, retrieval-augmented generation (RAG) leveraging external knowledge bases (KBs), referred to as KB-VQA, emerges as a promising approach. Nevertheless, conventional unimodal retrieval techniques, which translate images into textual descriptions, often result in the loss of critical visual details. This study presents fine-grained knowledge units, which merge textual snippets with entity images stored in vector databases. Furthermore, we introduce a knowledge unit retrieval-augmented generation framework (KU-RAG) that integrates fine-grained retrieval with MLLMs. The proposed KU-RAG framework ensures precise retrieval of relevant knowledge and enhances reasoning capabilities through a knowledge correction chain. Experimental findings demonstrate that our approach significantly boosts the performance of leading KB-VQA methods, achieving improvements of up to 10%.
@@
 Zhengxuan Zhang1, Yin Wu1, Yuyu Luo1,2, Nan Tang1,2thanks:   Nan Tang is the corresponding author 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology {zzhang393, ywu450}@connect.hkust-gz.edu.cn {yuyuluo, nantang}@hkust-gz.edu.cn
@@ 1 Introduction
Figure 1: Sample VQA Solutinon with MLLMs: (a) Direct answer without additional knowledge. (b) Single-modality coarse-grained RAG (KB-RAG). (c) Our proposal KU-RAG.
@@
Recently, the emergence of multimodal large language models (MLLMs), such as GPT-4 (Achiam et al., 2023) and LLaVA (Liu et al., 2023a), has introduced new possibilities for VQA. Unlike previous methods, MLLMs serve not only as powerful reasoning engines but also as vast knowledge repositories, with information learned from world knowledge during pretraining (Wang et al., 2024; Liu et al., 2023b). This dual capability enables more nuanced answers. However, the knowledge acquired during training is general and (maybe outdated) world knowledge, limiting the model’s ability to respond to domain-specific and update-to-date queries. As shown in Figure 1(a), when using GPT-4 to ask a question about the bridge in the image, it fails to provide an answer due to a lack of relevant knowledge and LLaVA even hallucinated and provided a “false” answer.
At this point, it becomes necessary to employ KB-VQA approachs, by retrieving information from a database – a process also known as Retrieval-Augmented Generation (RAG) in the context of LLMs (Fan et al., 2024). This typically involves converting images into captions and then performing passage-level retrieval combined with the query. However, this method struggles to handle fine-grained information for question answering, and during the image-to-text modality conversion process, some visual details are inevitably lost. As shown in Figure 1(b), a unimodal, coarse-grained approach fails to retrieve the relevant knowledge.
@@
Following this approach, we propose a “Knowledge Unit” component to bridge the query and specific knowledge. Specifically, we propose a Knowledge Unit Retrieval-Augmented Generation (KU-RAG) method, which is a multimodal, fine-grained, zero-shot retrieval approach covering both data storage and retrieval. As shown in the figure, our method matches the image from the question with images in the database, identifying the relevant knowledge (e.g., “Karnin Lift Bridge”) and subsequently retrieving fine-grained knowledge chunks related to it to answer the question more effectively. Finally, we designed a Knowledge Correction Chain (KCC) to assist in answer generation. The KCC integrates retrieved information into the MLLM’s reasoning process and verifies the accuracy of the knowledge generated by MLLMs.
Our contributions are summarized as follows.
  • We introduce knowledge units that consist of fine-grained multimodal data fragments (e.g., text fragments, entity images, and so on) from external knowledge bases. We store these knowledge units in vector databases, in order to effectively retrieve relevant knowledge for a given visual-based query.
@@
  • The experimental results on various KB-VQA benchmarks demonstrate that our method effectively enhances the performance of MLLM on this task.
2 Related Work
2.1 Knowledge-based Visual Question Answering
Knowledge-based Visual Question Answering (KB-VQA) uses external knowledge like Wikipedia to assist in answering image-related questions Wu et al. (2022). Early methods used a retriever-reader framework, retrieving relevant text to predict answers, but struggled with fine-grained questions about visual entities (Hu et al., 2023a), events (Yang et al., 2023), and visual information (Chen et al., 2023).
With the rise of LLMs, some methods combined implicit LLM knowledge with database retrieval, converting images into tags or captions and using GPT for knowledge retrieval (Gui et al., 2021; Chen et al., 2024; Wu et al., 2025). However, a gap exists between queries and LLM knowledge. Hu et al. (2023b) proposed a prompt-guided image captioning method to adjust captions based on queries. Although some methods reduce information loss through visual features (Salaberria et al., 2023) or enriched prompts (Shao et al., 2023), they don’t fully resolve the issue. Our approach integrates multimodal information for both retrieval and reasoning, with the LLM acting as both a knowledge base and a reasoner.
2.2 Multimodal Retrieval-augmented Generation
Multimodal Retrieval-augmented Generation (RAG) improves LLMs by retrieving relevant external documents to address domain-specific knowledge and avoid hallucinations (Gao et al., 2023). The process retrieves documents and generates answers by integrating the information, similar to KB-VQA’s retrieval. RAG has evolved with models like GRAG, which improves retrieval relevance (Hu et al., 2024), FiD-RAG, which fuses multiple documents during generation (Izacard and Grave, 2020), and DPR-RAG, which uses dense retrieval to find relevant fragments quickly (Karpukhin et al., 2020).
Unlike the aforementioned methods, our goal is to build a multimodal, fine-grained model that combines the internal knowledge of LLMs to generate high-quality answers in zero-shot scenarios.
3 Our Methodology: VQA with Knowledge Unit RAG and MLLMs
In this section, we will provide a detailed explanation of the Knowledge Unit Retrieval-Augmented Generation (KU-RAG) method. We construct a novel framework based on KU-RAG for KB-VQA, as illustrated in Figure 2. We will introduce the KB-VQA task in Section 3.2, then present our proposed “knowledge unit” in Section 3.2, followed by a complete overview of KU-RAG in Section 3.3.
Figure 2: The KU-RAG Framework
@@ 3.1 Task Definition
Visual Question Answering (VQA): Given a question $Q$, which consists of an image $Q_{i}$ and a textual question $Q_{t}$ related to the content of the image, the task of VQA is to generate an answer $A$ based on the information available in the image and the text. In this setup, the system aims to understand both the visual and textual aspects of the input and provide a relevant response.
Knowledge-Based Visual Question Answering (KB-VQA): In KB-VQA, the goal extends beyond the basic VQA task by incorporating external knowledge $K$ stored in knowledge bases (KBs) to answer the question. This external knowledge, which can be categorized as either image knowledge $K_{i}$ or text knowledge $K_{t}$, is retrieved based on the question $Q$ and is used to generate a more informed and accurate answer $A$.
@@ 3.2 Knowledge Unit
3.2.1 Definition
We introduce a new structure called Knowledge Unit (KU). Each KU serves as a knowledge carrier or object generated in combination with the query, such as entities, events, rules, topics, etc., designed to bridge the gap between the query and the database during the actual question-answering process.
@@ 3.2.1 Definition
In the KB-VQA task, image-image or name-name matching is typically used to determine which piece of knowledge a given image belongs to. Hence, we encapsulate $K_{i}$ and $K_{n}$ into the ‘Matching End’ to link the query and the KU. The purpose of KB-VQA often involves querying the knowledge behind an image, so we refer to $K_{t}$ as the ‘Detail End’.
3.2.2 Construction
Knowledge Predefinition.
@@ 3.2.2 Construction
Firstly, we should determine how to extract the knowledge unit with the application scenario and consider the query and database. For example, in an object recognition QA system, different entities can serve as knowledge units; in an event query system, different events can serve as knowledge units; in a corporate rules and regulations query system, a rule-based knowledge unit should be constructed. It is important to note that knowledge units do not necessarily need to be atomic or as fine-grained as possible. Its division should be determined based on the granularity of the data in the query and the database. For instance, a general animal knowledge QA system may only require the general species of an animal (e.g., “cat”), whereas a specific species QA system may require the specific species name (e.g., “Persian cat”).
Knowledge Segment.
@@
Since subsequent steps involve the storage of the knowledge unit, storing textual knowledge $K_{t}$ within the detail end at the document level, which is a coarse-grained storage method, is highly detrimental to knowledge retrieval (Chen et al., 2021). Additionally, during reasoning, this can be limited by the LLM’s maximum token capacity, leading to incomplete information. Therefore, we first need to segment the raw data, breaking it down into finer-grained units.
@@
  • Textual data: Similar to general RAG methods, we next segment all text passage $P=(p_{1},p_{2},\dots,p_{n})$ in a knowledge base to obtain the smallest retrieval units. Each passage $P$ contains $n$ sentences, i.e., $P_{k}=(s_{1},s_{2},\dots,s_{n})$. Considering the importance of knowledge coherence in the KB-VQA task, we adopt a combination of sentence splitting and maximum token limit. In each chunk, as many sentences as possible are retained without exceeding the maximum token limit, and the remaining sentences are assigned to the next chunk. That is, $C=(c_{1},c_{2},\dots c_{i})$ , where $c_{j}=(s_{1},s_{2},\dots s_{j})$. $i$ represents the $i$-th chunk, and $j$ represents the $j$-th sentence in the chunk.
@@ 3.2.3 Storage
$V_{c_{i}}=\text{Encoder}(c_{i}),D_{c}=(V_{c1},V_{c2},\dots,V_{cn})$ (2)
@@
$V_{i_{j}}=\text{Encoder}(i_{j}),D_{i}=(V_{i_{1}},V_{i_{2}},\dots,V_{i_{n}})$ (3)
Figure 3: The construction of knowledge units (KU) with entity and event. The entity list comes from the work of  Hu et al. (2023a), while the events are sourced from E-VQA (Yang et al., 2023).
3.3 Knowledge Unit Retrieval Framework
In this section, we will introduce our knowledge unit retrieval framework and detail how to achieve knowledge retrieval through knowledge unit and apply it to the KB-VQA task.
As shown in Figure 2, our framework is divided into three modules: Knowledge Unit Construction, Knowledge Retrieval, and MLLM Reasoning & Generation. The knowledge unit construction module mainly transforms raw knowledge into knowledge units and stores them in the database, as illustrate in Section 3.2. The knowledge retrieval module processes the original query, matches it with the corresponding knowledge units, finds the relevant knowledge, and integrates it into a structured, MLLM-readable passage. Finally, combining the original question and the retrieved knowledge, the MLLM Reasoning & Generation module analyzes and generates the answer.
3.3.1 Query Processing
For the user’s input query $Q$ , it is first necessary to preprocess and rewrite it to reduce interference during the retrieval process. To find the region in the image related to the question, we propose a query-aware instance segmentation method. Specifically, we first use YOLO (Redmon et al., 2016) to perform instance segmentation on the image, obtaining n segmented instance objects $O=(o_{1},o_{2},\dots,o_{n})$ We then encode these instances using Long-CLIP, resulting in corresponding vectors $V_{o}=(v_{o_{1}},v_{o_{2}},\dots,v_{o_{n}})$:
@@ 3.3.1 Query Processing
$V_{q_{t}}=\text{Encoder}(Q_{t}),S_{n}=\text{Sim}(V_{q_{t}},V_{o_{i}})$ (5)
Here, $S_{n}=(s_{1},s_{2},\dots,s_{n})$ represents the similarity values corresponding to each O . We select the object o with the highest similarity that exceeds a threshold $\gamma$ for subsequent retrieval, with its vector denoted as $V_{q_{i}}$ Of course, sometimes the query may not only be related to one object but also to areas outside the objects or the entire scene in the image. Therefore, if no object meets the criteria or multiple objects meet the criteria, we will encode the entire image and use this encoding for subsequent retrieval:
@@
3.3.2 Knowledge Unit Matching
@@
$KU=\text{Matching}(V_{q_{i}}),KU^{\prime}=\text{top-k}(KU)$ (7)
@@
3.3.3 MLLM Reasoning and Generation
@@
Modal Aligning and Fusing: First, based on the retrieval results $C^{\prime}$, we find the corresponding knowledge unit $KU^{\prime}$ for each chunk and combine its matching end information to form an image with the structure ‘[Image][[Name][Chunk Text]]’, where the image corresponding to the $i$-th chunk is denoted as $I^{\prime}_{i}$ Notably, if multiple chunks correspond to the same image, to enhance the connection between knowledge and improve processing efficiency, we merge the texts of these chunks into a single image in the format ‘[Image][[Name][Chunk $Text_{1}$]…[Chunk $Text_{n}$]]’.
Images Stitching: Next, we stitch all the images obtained in the previous step to generate a multimodal passage with both image and text information. This multimodal passage $MP$ is as:
@@
Knowledge Correction Chain: At this stage, a key challenge is effectively managing the relationships among the ‘information in the query’, ‘the knowledge retrieved’, and ‘the inherent knowledge of the MLLM’, as well as ensuring a fine-grained correspondence between text and images.
In our experiments, we found that when combining the retrieved knowledge with the question for the MLLM to answer, the model tended to prioritize the retrieved information while neglecting its own knowledge. We also attempted to use guiding prompts, such as “Based on your own knowledge first…” and “Focus on the first image and ignore other image…” to encourage the MLLM to consider its own knowledge before referring to the retrieved information, but the results were unsatisfactory (as demonstrated in Section  4).
To address this issue, we design a Knowledge Correction Chain (KCC) that guides MLLMs in reasoning through multi-turn dialogue and reading comprehension. In details, we first input the question $Q$ to MLLM to obtain the original answer $A_{0}$:
@@
The purpose of this step is to obtain the pure knowledge of the MLLM regarding the query, without being influenced by the retrieved information mentioned above. Finally, we input the passage $MP$ into the MLLM with a knowledge correction aware (KC-aware) prompt and get the final answer $A$:
@@
In this way, we can fully utilize multimodal information and handle the fine-grained correspondences between them, enhancing the MLLMs’ ability to reason and answer questions in complex scenarios. In addition, we have designed a method for managing KU. More details can be referred to in Appendix  A.

### v2

@@ Abstract
Visual Question Answering (VQA) focuses on providing answers to natural language questions by utilizing information from images. Although cutting-edge multimodal large language models (MLLMs) such as GPT-4o achieve strong performance on VQA tasks, they frequently fall short in accessing domain-specific or the latest knowledge. To mitigate this issue, retrieval-augmented generation (RAG) leveraging external knowledge bases (KBs), referred to as KB-VQA, emerges as a promising approach. Nevertheless, conventional unimodal retrieval techniques, which translate images into textual descriptions, often result in the loss of critical visual details. This study presents fine-grained knowledge units, which merge textual snippets with entity images stored in vector databases. Furthermore, we introduce a knowledge unit retrieval-augmented generation framework (KU-RAG) that integrates fine-grained retrieval with MLLMs. The proposed KU-RAG framework ensures precise retrieval of relevant knowledge and enhances reasoning capabilities through a knowledge correction chain. Experimental findings demonstrate that our approach significantly boosts the performance of leading KB-VQA methods, achieving an average improvement of approximately 3% and up to 11% in the best case.
@@
 Zhengxuan Zhang1thanks:   Zhengxuan Zhang and Yin Wu contributed equally to this work., Yin Wu [^1] footnotemark: 1, Yuyu Luo1,2, Nan Tang1,2thanks:   Nan Tang is the corresponding author 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology {zzhang393, ywu450}@connect.hkust-gz.edu.cn {yuyuluo, nantang}@hkust-gz.edu.cn
@@ 1 Introduction
Figure 1: Sample VQA Solution with MLLMs: (a) Direct answer without additional knowledge. (b) Single-modality coarse-grained RAG. (c) Our proposal KU-RAG.
@@
VQA with MLLMs. Recently, the emergence of multimodal large language models (MLLMs), such as GPT-4 (Achiam et al., 2023) and LLaVA (Liu et al., 2023a), has introduced new possibilities for VQA. Unlike previous methods, MLLMs serve not only as powerful reasoning engines but also as vast knowledge repositories, with information learned from world knowledge during pretraining (Wang et al., 2024; Liu et al., 2023b). This dual capability enables more nuanced answers. However, the knowledge acquired during training is general and (maybe outdated) world knowledge, limiting the model’s ability to respond to domain-specific and update-to-date queries. As shown in Figure 1(a), when using GPT-4 to ask a question about the bridge in the image, it fails to provide an answer due to a lack of relevant knowledge, and LLaVA even hallucinated and provided a “false” answer.
VQA with RAG and MLLMs. At this point, it becomes necessary to employ KB-VQA methods by retrieving information from a database – a process also known as Retrieval-Augmented Generation (RAG) in the context of LLMs (Fan et al., 2024). This typically involves converting images into captions and then performing passage-level retrieval combined with the query. However, this method struggles to handle fine-grained information for question answering, and during the image-to-text modality conversion process, some visual details are inevitably lost. As shown in Figure 1(b), a unimodal, coarse-grained approach fails to retrieve the relevant knowledge.
@@
Our Proposal: VQA with MLLMs and Fine-Grained Structured Knowledge. Following this approach, we propose a “Knowledge Unit” (KU) component to bridge the query and specific knowledge. Unlike traditional RAG methods that primarily focus on optimizing retrieval strategies, our approach centers on designing a structured knowledge representation that naturally connects queries with relevant knowledge. As shown in Fig. 1 (c), the KU consists of the picture and the description of the Karnin Lift Bridge.
Based on it, we propose a Knowledge Unit Retrieval-Augmented Generation (KU-RAG) method, which is a multimodal, fine-grained, zero-shot retrieval approach covering both data storage and retrieval. As shown in the figure, our method matches the image from the question with images in the database, identifying the relevant knowledge (e.g., “Karnin Lift Bridge”). Instead of merely refining retrieval mechanisms, we emphasize the structured management of knowledge units, where the process of organizing and managing knowledge itself optimizes retrieval quality. Finally, we designed a Knowledge Correction Chain (KCC) to assist in answer generation. The KCC integrates retrieved information into the MLLM’s reasoning process and verifies the accuracy of the knowledge generated by MLLMs.
Contributions. Our key contributions are summarized as follows:
  • We introduce the concept of Knowledge Units (KUs), which structure fine-grained multimodal knowledge for efficient retrieval in database systems. By focusing on the structured management of KUs, we enhance retrieval quality beyond just refining retrieval mechanisms.
@@
  • We evaluate our approach on multiple KB-VQA benchmarks, demonstrating its effectiveness in improving knowledge retrieval and reasoning capabilities in VQA.
2 Preliminary
@@ 2.1 Task Definition
Visual Question Answering (VQA). Given a question $Q$, which consists of an image $Q_{i}$ and a textual question $Q_{t}$ related to the content of the image, the task of VQA is to generate an answer $A$ based on the information available in the image and the text. In this setup, the system aims to understand both the visual and textual aspects of the input and provide a relevant response.
Knowledge-Based Visual Question Answering (KB-VQA). In KB-VQA, the goal extends beyond the basic VQA task by incorporating external knowledge $K$ stored in knowledge bases (KBs) to answer the question. This external knowledge, which can be categorized as either image knowledge $K_{i}$ or text knowledge $K_{t}$, is retrieved based on the question $Q$ and is used to generate a more informed and accurate answer $A$.
@@ 2.2 Knowledge Unit
For KB-VQA, the core challenge lies in accurately locating relevant knowledge. As shown in Figure 2, general coarse-grained retrieval methods, such as image-to-image or caption-based text retrieval, often introduce significant noise or cause loss of visual information, making it difficult to retrieve the correct knowledge. Fine-grained retrieval alleviates some of this noise by focusing on predefined text entities or visual objects, but still fails to preserve complete visual context and cannot effectively integrate textual knowledge at the entity level.
To address this, We introduce a new structure called Knowledge Unit (KU). Each KU serves as a knowledge carrier or object generated in combination with the query, such as entities, events, rules, topics, etc., designed to bridge the gap between the query and the database during the actual question-answering process. Figure 3 illustrates examples of KUs constructed as entities and events.
@@
Figure 2: Illustration of general coarse-grained retrieval, fine-grained retrieval and knowledge unit retrieval for VQA.
@@
In the KB-VQA task, image-image or name-name matching is typically used to determine which piece of knowledge a given image belongs to. Hence, we encapsulate $K_{i}$ and $K_{n}$ into the “Matching End” to link the query and the KU. The purpose of KB-VQA often involves querying the knowledge behind an image, so we refer to $K_{t}$ as the “Detail End”. As shown in the third figure of Fig. 2, we take the Karnin Lift Bridge as an example to construct a KU. During actual operation, the user provides an image $Q_{i}$, which is matched through the Matching End against existing images. The system successfully identifies the KU corresponding to the Karnin Lift Bridge. Subsequently, background knowledge related to this KU is retrieved via its Detail End, and the relevant knowledge fragment is located with the help of the query $Q_{t}$.
Our objective is to manage knowledge effectively through the structured organization and manipulation of KUs, thereby enabling seamless bridging between user queries and knowledge. To achieve this, there are two main challenges:
  • How to classify and extract knowledge from a large amount of unannotated text to construct structured KUs (i.e., the KU construction problem);
  • How to support flexible KU operations such as insertion, deletion, and updating (i.e., the KU operation problem).
We present our solutions to these challenges in Section 3.
3 Knowledge Unit Management
3.1 Construction
Figure 3: The construction of knowledge units with entity and event. The entity list comes from the work of  Hu et al. (2023a), while the events are sourced from E-VQA (Yang et al., 2023).
Figure 4: Overview of KU-RAG Framework
@@ 3.1 Construction
Knowledge Predefinition. Firstly, we should determine how to extract the knowledge unit with the application scenario and consider the query and database. For example, in an object recognition QA system, different entities can serve as knowledge units; in an event query system, different events can serve as knowledge units; in a corporate rules and regulations query system, a rule-based knowledge unit should be constructed. It is important to note that knowledge units do not necessarily need to be atomic or as fine-grained as possible. Its division should be determined based on the granularity of the data in the query and the database. For instance, a general animal knowledge QA system may only require the general species of an animal (e.g., “cat”), whereas a specific species QA system may require the specific species name (e.g., “Persian cat”).
@@
Knowledge Segment. Since subsequent steps involve the storage of the knowledge unit, storing textual knowledge $K_{t}$ within the detail end at the document level, which is a coarse-grained storage method, is highly detrimental to knowledge retrieval (Chen et al., 2021). Additionally, during reasoning, this can be limited by the LLM’s maximum token capacity, leading to incomplete information. Therefore, we first need to segment the raw data, breaking it down into finer-grained units.
@@
  • Textual data: Similar to general RAG methods, we next segment all text passage $P=(p_{1},p_{2},\dots,p_{n})$ in a knowledge base to obtain the smallest retrieval units. Each passage $P$ contains $n$ sentences, i.e., $P_{k}=(s_{1},s_{2},\dots,s_{n})$. Considering the importance of knowledge coherence in the KB-VQA task, we adopt a combination of sentence splitting and maximum token limit. In each chunk, as many sentences as possible are retained without exceeding the maximum token limit, and the remaining sentences are assigned to the next chunk. That is, $C=(c_{1},c_{2},\dots c_{i})$ , where $c_{j}=(s_{1},s_{2},\dots s_{j})$, $i$ represents the $i$-th chunk, and $j$ represents the $j$-th sentence in the chunk.
@@
$V_{c_{i}}=\text{Encoder}(c_{i})$ (2)
$D_{c}=(V_{c1},V_{c2},\dots,V_{cn})$ (3)
@@
$V_{i_{j}}=\text{Encoder}(i_{j})$ (4)
$D_{i}=(V_{i_{1}},V_{i_{2}},\dots,V_{i_{n}})$ (5)
3.2 Database-level Operation
Knowledge Addition. When new knowledge is introduced, we directly chunk the corresponding text. The encoded text vectors are then stored in the existing text vector database $D_{c}$. Similarly, for images within the knowledge base, we encode them and store the resulting vectors in the image vector database $D_{i}$.
Knowledge Unit Operation. After performing operations on the raw data, it is necessary to consider whether corresponding actions need to be taken for the knowledge unit.
4 Knowledge Unit Retrieval Framework
In this section, we will introduce our knowledge unit retrieval framework and detail how to achieve knowledge retrieval through knowledge units and apply it to the KB-VQA task.
As shown in Figure 4, our framework is divided into three modules: Knowledge Unit Construction, Knowledge Retrieval, and MLLM Reasoning & Generation. The knowledge unit construction module mainly transforms raw knowledge into knowledge units and stores them in the database, as illustrated in Section 2.2. The knowledge retrieval module processes the original query, matches it with the corresponding knowledge units, finds the relevant knowledge, and integrates it into a structured, MLLM-readable passage. Finally, combining the original question and the retrieved knowledge, the MLLM Reasoning & Generation module analyzes and generates the answer.
4.1 Query Processing
For the user’s input query $Q$ , it is first necessary to preprocess and rewrite it to reduce interference during the retrieval process. To find the region in the image related to the question, we propose a query-aware instance segmentation method. Specifically, we first use YOLO (Redmon et al., 2016) to perform instance segmentation on the image, obtaining n segmented instance objects $O=(o_{1},o_{2},\dots,o_{n})$. We then encode these instances using Long-CLIP, resulting in corresponding vectors $V_{o}=(v_{o_{1}},v_{o_{2}},\dots,v_{o_{n}})$:
@@ 4.1 Query Processing
$V_{q_{t}}=\text{Encoder}(Q_{t})$ (7)
$S_{n}=\text{Sim}(V_{q_{t}},V_{o_{i}})$ (8)
Here, $S_{n}=(s_{1},s_{2},\dots,s_{n})$ represents the similarity values corresponding to each O . We select the object o with the highest similarity that exceeds a threshold $\gamma$ for subsequent retrieval, with its vector denoted as $V_{q_{i}}$. Of course, sometimes the query may not only be related to one object but also to areas outside the objects or the entire scene in the image. Therefore, if no object meets the criteria or multiple objects meet the criteria, we will encode the entire image and use this encoding for subsequent retrieval:
@@
4.2 Knowledge Unit Matching
@@
$KU=\text{Matching}(V_{q_{i}})$ (10)
$KU^{\prime}=\text{top-k}(KU)$ (11)
@@
5 MLLM Reasoning and Generation
@@
Modal Aligning and Fusing. First, based on the retrieval results $C^{\prime}$, we find the corresponding knowledge unit $KU^{\prime}$ for each chunk and combine its matching end information to form an image with the structure ‘[Image][[Name][Chunk Text]]’, where the image corresponding to the $i$-th chunk is denoted as $I^{\prime}_{i}$. Notably, if multiple chunks correspond to the same image, to enhance the connection between knowledge and improve processing efficiency, we merge the texts of these chunks into a single image in the format ‘[Image][[Name][Chunk $Text_{1}$]…[Chunk $Text_{n}$]]’.
Images Stitching. Next, we stitch all the images obtained in the previous step to generate a multimodal passage with both image and text information. This multimodal passage $MP$ is as:
@@
Knowledge Correction Chain. At this stage, a key challenge is effectively managing the relationships among the “information in the query”, “the knowledge retrieved”, and “the inherent knowledge of the MLLM”, as well as ensuring a fine-grained correspondence between text and images.
In our experiments, we found that when combining the retrieved knowledge with the question for the MLLM to answer, the model tended to prioritize the retrieved information while neglecting its own knowledge. We also attempted to use guiding prompts, such as “Based on your own knowledge first…” and “Focus on the first image and ignore other image…” to encourage the MLLM to consider its own knowledge before referring to the retrieved information, but the results were unsatisfactory (as demonstrated in Section  6).
To address this issue, we design a Knowledge Correction Chain (KCC) that guides MLLMs in reasoning through multi-turn dialogue and reading comprehension. In detail, we first input the question $Q$ to MLLM to obtain the original answer $A_{0}$:
@@
The purpose of this step is to obtain the pure knowledge of the MLLM regarding the query without being influenced by the retrieved information mentioned above. Finally, we input the passage $MP$ into the MLLM with a knowledge correction aware (KC-aware) prompt and get the final answer $A$:
@@
In this way, we can fully utilize multimodal information and handle the fine-grained correspondences between them, enhancing the MLLMs’ ability to reason and answer questions in complex scenarios.